# Importance of accounting for student identities and intersectionality for creating equitable and inclusive physics learning environments

### v1

@@ Abstract
This research focuses on the experiences of seven undergraduate women who were majoring in physics at a small liberal arts college. In the semi-structured, empathetic interviews we conducted, the women discuss how they decided to major in physics, their interactions with their peers and instructors, who has supported them during their physics trajectory, and suggestions that would improve their experiences in physics. We use Standpoint Theory and focus on the experiences of undergraduate women to get a holistic perspective of how they became interested in physics, how they have been supported in their physics journey as well as identify any challenges that they faced in their undergraduate physics program due to their identity. Using synergistic frameworks such as the domains of power and the Holistic Ecosystem for Learning Physics in an Inclusive and Equitable Environment (HELPIEE), we analyze how those in the position of power, e.g., instructors, can play important roles in establishing and maintaining safe, equitable, and inclusive environments for students, which is especially important for historically marginalized students such as women and ethnic and racial minority students in physics. Using these frameworks, we compare the experiences of these women with those in Johnson’s research at a small college and our prior investigation at a large research university. We also discuss the suggestions provided by the undergraduate women to implement in the future to support current and future undergraduate women in physics and astronomy. Their suggestions are separated as personal advice for peers and suggestions for physics instructors.
preprint: APS/123-QED
I Introduction, Framework and Goal
Many groups, e.g., women and ethnic and racial minorities (ERM), have historically been excluded from physics and improving their representation and experiences has been an important focus for researchers for the past few decades. Within the last decade, the percentage of women who received physics bachelor’s degrees in the US has ranged from 20% to 25% [1, 2, 3]. Reports from 2020 shows that of physics bachelor’s degrees, 3-4% are awarded to Black students and 12% are awarded to Latinx students [4, 5].
@@ I Introduction, Framework and Goal
At many institutions, physics is still tied to its historical origins in regard to its culture. For example, physics has the worst stereotypes among the natural sciences regarding who belongs in it, who can excel in it and what a physicist looks like [24, 25]. As a result of these stereotypes, students are negatively affected [21], including in their performance [26, 27, 28]. In addition, the demographics of physics departments may continue to reinforce stereotypes, and historically marginalized students, who don’t have role models, continue to be constantly reminded that they do not necessarily fit in.
@@
Previous studies have shown gender disparities in performance as well as psychological factors in physics persisting throughout high school and college. Such disparities include women’s lower perceived recognition compared to men as a “physics person” from peers and instructors [29, 30, 31, 32, 33, 34, 35, 36], which can hinder their academic performance [37] and influence female students’ decisions to leave the field [38, 6, 7, 39]. In particular, if women believe that their instructors or peers do not see them as being capable of excelling in physics, it impacts their own beliefs about whether they can excel in physics [40, 41, 34, 42]. Moreover, lower self-efficacy and sense of belonging can be detrimental even to those students who are performing well. Prior research also shows that women drop out of STEM disciplines with significantly higher grades than men [43]. Prior studies have also found that sense of belonging and self-efficacy in physics are closely intertwined [44, 45, 46]. Because improving self-efficacy is a multifaceted process [47], creating an equitable and inclusive learning environment that increases underrepresented students’ sense of belonging may help increase their self-efficacy and improve their retention [48, 49, 50, 51, 52, 53, 54, 55].
@@
In order to do this, physics instructors must be committed to these issues in order to establish and maintain a safe and equitable physics learning environment. They also have the power to do so as they are responsible for setting the tone of how students are treated and establish classroom norms. Another important aspect to emphasize is that in many physics programs, men make up the majority or both physics majors and physics faculty. Hence, they have both power and influence to make change. We also note that physics culture has traditionally been masculine, as it still is at many universities, which continues to perpetuate inequities and stereotypes for traditionally marginalized students. The physics culture proclaims and reinforces ideas that only brilliant people can do physics and since men are stereotypically associated with brilliance [56], they are the only ones who can do physics and belong in physics [8, 57, 28, 58, 59, 60, 61, 62, 26, 63, 64, 65]. This masculine culture disadvantages women and they become isolated from lack of role models and a community that could provide support to them [66, 67, 65, 68].
@@
Thus, is it important to investigate the experiences of underrepresented students in physics, such as women as in our study, as they are the ones who can shed light on their interactions with others in physics, and how the physics culture impacts them. Qualitative studies such as this investigation can be valuable to learn about inequities directly from students who may be experiencing it. Through interviews, we can understand undergraduate women’s perspectives in their own voices. Previous studies in other contexts that utilize student interviews revealed that women in physics and astronomy experience hostile environments within physics departments [19, 69, 21, 60, 22, 23].
@@
We choose Standpoint theory as our theoretical framework to help frame this study in which we put undergraduate women in physics at the forefront [70, 71]. Standpoint theory is a critical theory that focuses on the relationship between the production of knowledge and acts of power [72]. It is related to other critical theories in that it centers around the standpoint or voices of the underrepresented groups that do not have the same privilege as the dominant group in order to gain a clearer understanding of their struggles. Thus, using this framework, we put emphasis on the experiences of undergraduate women in physics to understand what physics departments and instructors can do to improve the physics culture such that they feel safe and supported [73].
In this study, we use qualitative interview data to understand undergraduate women’s trajectory in physics (including what or who got them interested in physics), interactions with their peers and instructors, who supported them in physics and challenges they face, and their suggestions to improve their experiences. We compare these findings to Johnson’s study at a small college and a study we conducted several years ago at a large research university to analyze if there were major differences in women’s interactions with male peers and faculty across these different types of institutions [74, 23]. In the previous study conducted by Johnson at a small college, she analyzed women of color’s experiences using Collins’ domains of power framework. We compare undergraduate women’s experiences at these different institutions using a similar analysis employing the domains of power framework after discussing our results at the small liberal arts college presented here. We also make note of the fact that the domains of power framework is closely aligned with the Holistic Ecosystem for Learning Physics in an Inclusive and Equitable Environment (HELPIEE) framework which focuses on the role of those in the position of power in creating an equitable and inclusive learning environment.
I.1 Domains of Power and HELPIEE Frameworks
@@
Collins introduced four domains that are necessary to understand how power is organized in a particular context [75]. She argued that the four domains we need to consider are: interpersonal, cultural, structural, and disciplinary. The interpersonal domain is where power is expressed between individuals. The cultural domain is where group values are expressed, maintained or challenged. The disciplinary domain is how rules are enforced and for whom. The structural domain is how power is organized in various structures. Johnson applied this framework in the context of physics learning environments at a small college where women of color felt supported and were excited about learning physics. With the domains of power framework, it is the instructors who have the power over shaping the narrative for each of these domains (which interact with each other). In the context of student experiences in learning physics, the interpersonal domain refers to how students communicate with each other and how students and faculty interact; the cultural domain refers to the physics culture; the structural domain refers to the structure of a physics learning environment; and the disciplinary domain refers, e.g., to how instructors discipline students in the physics courses and in the physics learning environments in general, if their conduct does not conform do the expected norms.
@@
The Holistic Ecosystem for Learning Physics in an Inclusive and Equitable Environment (HELPIEE) framework posits that those in the position of power, e.g., instructors in physics classrooms, have the power to help all students feel supported by carefully taking into account students’ characteristics and implementing effective pedagogies in an equitable and inclusive learning environment [76]. Within this framework, if students from different demographic groups in a course do not have similar positive experiences and feelings of being supported, the learning environments are not equitable because those in the position of power did not provide adequate support to level the playing field.
Both of these frameworks are synergistic. From these two frameworks, we see that in the context of physics courses, it is the department and instructor’s responsibility to create an equitable learning environment. Together, these frameworks are useful for analyzing the findings of this study and comparing them with prior studies. Below, we describe our research questions.
I.2 Research Questions
  1. RQ1.
    What are the experiences of undergraduate women majoring in physics at a small liberal arts college?
  2. RQ2.
    How do the results of this study compare to other studies with similar themes but in different types of physics learning environments?
II Methodology
@@
We conducted semi-structured, empathetic interviews with 7 undergraduate women physics and astronomy majors at a small liberal arts college in the US. At this college, women are underrepresented in physics, but not to the extent that they are underrepresented at the national level [1], i.e., they make up about one-third of the physics majors. We note that there are 9 full-time faculty members and 4 visiting/adjunct faculty members in this physics and astronomy department. The number of physics majors that graduate each year varies from 10-20. In order to capture a wide range of experiences, we interviewed women who were at various points in their physics trajectory in college, e.g., second to fourth year students.
Each interview was about an hour long in duration and was recorded via Zoom. The interviews followed protocols set by the group prior to conducting the interviews. The protocol questions focused on these women’s experiences in physics, including their high school experiences and how their gender impacted their experiences. These women in our study volunteered to participate in the interviews.
We coded these interviews using hybrid coding (inductive and deductive) methods that emphasized these women’s experiences. Deductive codes are based on the interview questions while inductive codes were formed during the analysis [77]. The interview questions were centered around women’s experiences in physics, e.g., interactions with peers and instructors, why they majored in physics, etc. First, one of the researchers coded the interviews and both researchers engaged in several discussions related to the coding through which they came to an agreement on the codes. The codes were then distilled into several analytic themes. We note that in the excerpts, we removed most of the ”ums” and ”likes” to make it easier to follow their point of view.
II.1 Analytic Themes
@@ 6. AT6
II.2 Positionality
Both researchers identify as women of color. Specifically, one of the researchers identifies as a queer Chicana woman. The other is an Asian American woman. It is important to note these identities as they are also intersectional and were useful when coding and analyzing the intersectional identities of some of these undergraduate women in physics. Both have had several prior experiences conducting and analyzing qualitative studies focusing on equity and inclusion in physics learning environments.
II.3 Introduction of Women
Here we briefly introduce each woman in this study who was majoring in the physics and astronomy department at a small liberal arts college. All women were provided pseudonyms for anonymity. Cathy is a White fourth year student who is interested in computer science in addition to majoring in physics. Chloe is a White third year student. Esme is a Latina second year student. Francis is a White third year student who is majoring in math in addition to physics. Karina is a White fourth year student who is majoring in English in addition to physics. Paulina  [^1] We note that Esme and Paulina have different college experiences in regard to the aspects of their identities that have negatively impacted them and one reason for this could be the differences in their visual characteristics, i.e., only Paulina visually appears to be non-White. is a Latina third year student. Ruby is a White second year student. All of these women were brought up in the US, except Francis who grew up in the UK and came to the US to study at this college.
III Findings and Discussion
III.1 Pathway to Majoring in Physics

### v2

@@ Abstract
This research focuses on the experiences of seven undergraduate women who were majoring in physics in a medium-size physics department at a small liberal arts college. In the semi-structured, empathetic interviews we conducted, the women discussed how they decided to major in physics, their interactions with their peers and instructors, who supported them during their physics trajectory, and suggestions that would improve their experiences in physics. We used Standpoint theory and focused on the experiences of undergraduate women to get a holistic perspective of how they became interested in physics, how they have been supported in their physics journey as well as identify any challenges that they faced in their undergraduate physics program due to their identity. Using synergistic frameworks such as the Domains of Power and the Holistic Ecosystem for Learning Physics in an Inclusive and Equitable Environment (HELPIEE), we analyzed how those in the position of power, e.g., instructors, can play important roles in establishing and maintaining safe, equitable, and inclusive environments for students, which is especially important for historically marginalized students such as women and ethnic and racial minority students in physics. We also discuss the suggestions provided by the undergraduate women to implement in the future to support current and future undergraduate women in physics and astronomy. Their suggestions are separated as personal advice for peers and suggestions for physics instructors.
I Introduction
Many groups, e.g., women and ethnic and racial minorities (ERM), have historically been underrepresented from physics and improving their representation and experiences has been an important focus for researchers for the past few decades. Within the last decade, the percentage of women who received physics bachelor’s degrees in the US has ranged from 20% to 25% [1, 2, 3]. Reports from 2020 shows that of physics bachelor’s degrees, 3-4% are awarded to Black students and 12% are awarded to Latinx students [4, 5].
@@ I Introduction
At many institutions, physics is still tied to its historical origins in regard to its culture. For example, physics has stereotypes among the natural sciences regarding who belongs in it, who can excel in it and what a physicist looks like [24, 25]. Students are negatively affected as a result of these stereotypes [21], including in their performance [26, 27, 28]. In addition, the demographics of physics departments may continue to reinforce stereotypes, meaning historically marginalized students, who don’t have role models, continue to be constantly reminded that they do not necessarily fit in.
@@
Previous studies have shown gender disparities in performance as well as psychological factors in physics persisting throughout high school and college. Such disparities include women’s lower perceived recognition compared to men as a “physics person” from peers and instructors [29, 30, 31, 32, 33, 34, 35, 36], which can hinder their academic performance [37] and influence women’s decisions to leave the field [38, 6, 7, 39]. In particular, if women believe that their instructors or peers do not see them as being capable of excelling in physics, it can impact their own beliefs about whether they can excel in physics [40, 41, 34, 42]. Moreover, lower self-efficacy and sense of belonging can be detrimental even to those students who are performing well. Prior research also shows that women drop out of STEM disciplines with significantly higher grades than men [43]. Prior studies have also found that sense of belonging and self-efficacy in physics are closely intertwined [44, 45, 46]. Because improving self-efficacy is a multifaceted process [47], creating an equitable and inclusive learning environment that increases underrepresented students’ sense of belonging may help increase their self-efficacy and improve their retention [48, 49, 50, 51, 52, 53, 54, 55].
@@
We also note that physics culture has traditionally been masculine, as it still is at many universities, which continues to perpetuate inequities and stereotypes for traditionally marginalized students. The physics culture proclaims and reinforces ideas that only brilliant people can do physics and since men are stereotypically associated with brilliance [56], they are the only ones who can do physics and belong in physics [8, 57, 28, 58, 59, 60, 61, 62, 26, 63, 64, 65]. This masculine culture disadvantages women and they become isolated from lack of role models and a community that could provide support to them [66, 67, 65, 68].
@@
It is important to investigate the experiences of underrepresented students in physics, such as women as in our study, since they are the ones who can shed light on their interactions with others in physics, and how the physics culture impacts them. Qualitative studies such as this investigation can be valuable to learn about inequities directly from students who may be experiencing it. Previous studies involving large research universities in the US in other contexts that utilize student interviews revealed that women in physics and astronomy experience hostile environments within these physics departments [19, 69, 21, 60, 22, 23].
In the US, there are many small liberal arts colleges with physics departments of various sizes (with regard to the number of faculty members and physics majors), which may potentially have very different physics cultures even if they are all liberal arts colleges. Moreover, the physics culture and its impact on traditionally underrepresented students such as women at these small institutions with physics departments of different sizes can be different from that at the large research universities that have been investigated in most prior qualitative research studies from the introductory to graduate levels. We also note that apart from Johnson’s study [70] in a small physics department at a small liberal arts college, we are not aware of prior qualitative studies that focus on the experiences of women physics majors in physics learning environments at small liberal arts colleges with physics departments of any size in the US. Also, to our knowledge, Santana and Singh qualitative study is the only one focusing exclusively on the experiences of undergraduate women majoring in physics [23] at a large research university in the US.
Furthermore, Kanim and Cid have emphasized that physics education researchers should design and implement studies at a variety of institutions with physics departments of different sizes so that the research is representative to the broader physics community and benefits students in different types of institutions and physics departments [71]. This investigation is an important step towards including the experiences and perspective of women physics majors in a medium-size physics department at a small liberal arts college. This study also provides opportunity to compare women physics majors’ experiences in a medium-size physics department with a small physics department in Johnson’s study [70] and a large research university [23] in the US in prior qualitative studies to better understand the similarities and differences in women physics majors’ experiences in physics departments of different sizes at various institutions. Through interviews, we can understand undergraduate women’s perspectives and their experiences in physics learning environments in their own voices.
This research uses qualitative data to investigate the experiences of undergraduate women majoring in physics in a medium-size physics department at a small liberal arts college in the US. In particular, we use qualitative interview data to understand undergraduate women’s trajectory in physics (including what or who got them interested in physics), interactions with their peers and instructors, who supported them in physics and challenges they face, and their suggestions to improve their experiences in the physics learning environments in a medium-size physics department at a small liberal arts college in the US.
Also, although a detailed comparison is out of the scope of this paper, we summarize how the findings at the medium-size physics department (presented in detail here) compare to Johnson’s study in a small physics department at a small liberal arts college and a study we conducted at a large research university to analyze if there were major differences in women’s interactions with male peers and faculty across these different types of institutions [70, 23].
II Frameworks
II.1 Standpoint Theory
@@
We choose Standpoint theory as our theoretical framework to help frame this study in which we put undergraduate women majoring in physics at the forefront [72, 73]. Standpoint theory is a critical theory that focuses on the relationship between the production of knowledge and acts of power [74]. It is related to other critical theories in that it centers around the standpoint or voices of the underrepresented groups that do not have the same privilege as the dominant group in order to gain a clearer understanding of their struggles. Thus, using this framework, we put emphasis on the experiences of undergraduate women in physics to understand what physics departments and instructors can do to improve the physics culture such that they feel safe and supported [75]. It is also important to acknowledge that women are experts in their own experiences, thus these interviews can provide valuable insight about how they navigate through various situations with this specific lens. From these interviews, we also obtain a lot of valuable information to improve the current physics learning environment, such as if and how they feel supported or what things must be improved upon. For example, these types of experiences can be used to inform/educate those in the position of power about how the women perceive dominant groups, or to reflect upon and evaluate one’s department to restructure the physics program/support systems. Thus, using this framework, we chose to interview only undergraduate women in physics in this medium-size physics department at a small liberal arts college.
II.2 Domains of Power
@@
Collins introduced four interconnected domains that are necessary to understand how power is organized in a particular context [76]. She argued that the four domains we need to consider are: interpersonal, cultural, structural, and disciplinary. The interpersonal domain pertains to how power is expressed between individuals. The cultural domain pertains to how group values are expressed, maintained or challenged. The structural domain pertains to how power is organized in various structures. The disciplinary domain pertains to how rules are enforced and for whom. Johnson applied this framework in the context of physics learning environments in a small physics department at a small college where women of color felt supported and were excited about learning physics. Within the Domains of Power framework, it is those in the position of power, e.g., instructors, who have the power over shaping the narrative for each of these domains (which interact with each other) in the physics learning environments [70]. In the context of student experiences in learning physics, the interpersonal domain refers to how students communicate with each other and how students and faculty interact; the cultural domain refers to the physics culture; the structural domain refers to the structure of a physics learning environment; and the disciplinary domain refers, e.g., to how instructors discipline students in the physics courses and in the physics learning environments in general, if their conduct does not conform to the expected norms.
II.3 HELPIEE
@@
The HELPIEE framework posits that those in the position of power, e.g., instructors in physics classrooms, have the power to help all students feel supported by carefully taking into account students’ characteristics and implementing effective pedagogies in an equitable and inclusive learning environment [77]. Student learning of physics can be affected by students’ backgrounds including their internal and external characteristics and whether learning environments are equitable and inclusive and provide appropriate support to all students based upon their needs. Internal characteristics include prior preparation, physics self-efficacy, identity etc. and external characteristics include their other academic and non-academic commitments and support systems such as the availability of appropriate advising/mentoring etc. Figure 1 summarizes the HELPIEE framework in which accounting for students’ characteristics can help instructors intentionally focus on student needs and tailor how to adapt and implement pedagogies in their courses to strengthen both students’ “Offense” or “Defense” skills for learning physics. Within this framework, if students from different demographic groups in a course, e.g., those from underrepresented groups, do not have similar positive experiences and feelings of being supported, the learning environments are not equitable because those in the position of power did not provide adequate support to level the playing field.
Figure 1: Figure from [77], summarizing the HELPIEE framework.
Both the Domains of Power and HELPIEE frameworks are synergistic. For example, in the context of physics courses, it is the department and instructor’s responsibility to create an equitable and inclusive learning environment. Together, these synergistic frameworks are useful for analyzing the findings of this study and comparing them with prior studies. Standpoint theory aligns with the two by putting the emphasis on women’s experiences to reveal the types of interactions they have in physics learning environments, what the physics culture is like, and what instructors’ actions are (or aren’t) in creating equitable learning environments. Thus, Standpoint theory acts as an overarching lens to focus on and learn from undergraduate women majoring in physics, who share their experiences. We also note that Standpoint theory also helps in the empathetic interviews by allowing the interviewer to connect with the participants based on this shared identity.
III Methodology
III.1 Participants
@@
We conducted semi-structured, empathetic interviews with 7 undergraduate women physics and astronomy majors in a medium-size physics department at a small liberal arts college in the US. At this college, women are underrepresented in physics, but not to the extent that they are underrepresented at the national level [1], i.e., they make up about one-third of the physics majors. We note that there are 9 full-time faculty members and 4 visiting/adjunct faculty members in this physics and astronomy department. The number of physics majors that graduate each year varies from 10-20. In order to capture a wide range of experiences, we interviewed women who were at various points in their physics trajectory in college, e.g., second to fourth year students.
CS contacted a faculty member at this college who then sent an email to all 20 undergraduate women (across different years) who were majoring in physics at the time. The email described the scope of the study and asked for those interested in participating to reach out to CS. Each of the seven participants (approximately one third of the women in that undergraduate physics program) volunteered to participate without compensation. At the beginning of each interview, the participants provided consent to be audio-recorded and quoted in academic publications. Each interview was about an hour long in duration and was recorded via Zoom. The interviews followed protocols set by the group prior to conducting the interviews (see Table 2). The protocol questions focused on the women’s experiences related to physics, such as what their high school experiences were in STEM, what their interactions were like with college peers and instructors, and whether and how their gender may have impacted their experiences. The interviews were semi-structured and we followed up on what the women described in response to the questions that were already part of the protocol.
We briefly introduce each woman in this study who was majoring in the physics and astronomy department in a medium-size physics department at this small liberal arts college in Table 1. All women were provided pseudonyms for anonymity. Several of the women (Cathy, Francis, and Karina) were either double majoring or interested in other subjects, such as computer science, mathematics, or English. All of the women were brought up in the US, except Francis who grew up in Europe and came to the US to study at this college.
Table 1: Here we provide information about the women that were interviewed.
Pseudonym Ethnic / Racial Background Year in college
Cathy White 4th year
Chloe White 3rd year
Esme Latina 2nd year
Francis White 3rd year
Karina White 4th year
Paulina [^1] We note that Esme and Paulina have different perspectives regarding which of their marginalized identities was prominent in their negative experiences. One reason for this could be the differences in their visual characteristics, i.e., only Paulina visually appears to be non-White. Latina 3rd year
Ruby White 2nd year
III.2 Positionality
Both researchers identify as women of color. Specifically, LMS identifies as a queer Chicana woman and CS identifies as an Asian American woman. It is important to note these identities as they are also intersectional and were useful when coding and analyzing the intersectional identities of some of these undergraduate women in physics. Both have had several prior experiences conducting and analyzing qualitative studies focusing on equity and inclusion in physics learning environments.
III.3 Data Analysis
We coded these interviews using hybrid coding (inductive and deductive) methods that emphasized the women’s experiences [78]. Deductive codes are based on the interview questions while inductive codes were formed during the analysis [78]. For example, based on the interview questions (see Table 2), we anticipated that the women would provide suggestions to improve their experiences although we did not anticipate the women to describe such specific suggestions that could be separated as personal advice for other women and advice for instructors. Another example of an inductive code was the reflections women provided regarding how their identities other than gender influenced how they navigated physics learning environments. Again, these themes emerged and were not directly based on the protocol questions specifically focusing on these.
To begin the analysis, LMS first corrected all the interviews from the Zoom transcriptions by listening to the audio-recordings to ensure that no information was misinterpreted. Both researchers read all the transcripts and had many extensive discussions about them with each other. Then, after these discussions between the two researchers, LMS proceeded to code the interviews using fine grain, specific codes, starting with one woman at a time. During this round, LMS and CS again engaged in several discussions about the codes including definitions and clear examples from the data. Through many back and forth discussions and fine-tuning, they came to an agreement on the general codes extracted from the fine grain codes. The general codes were then distilled into subthemes, and corresponding analytic themes [78]. In general, several subthemes were part of the same analytic theme. The researchers discussed and justified these subthemes and analytic themes based on how broad they were in elucidating the women’s experiences (capturing both negative and positive interactions with peers and instructors).
Figure 2: Layout of the data analysis, adapted from Figure $1.$ in [78]
Figure 2 provides a layout for how the codes were distilled starting from fine grain codes from individual participant data, to analytic themes. Fine grain codes are very specific for each participant and are not included here. In particular, depending upon their experiences, the women might have had similar or different fine grain codes for a particular question, which were then grouped into general codes. Thus, general codes are broader than fine grained codes to include data from all of the women. Table LABEL:tab:examples provides general codes, their definitions, and examples from individual participant data. Next, the general codes were distilled into subthemes. As shown in Table LABEL:tab:examples, these subthemes are broader than the general codes. Also, for illustration purposes, Figure 3 shows a subset of general codes which we color coded for a corresponding subtheme agreed upon by both researchers. As shown in Table LABEL:tab:examples, we grouped similar subthemes under Analytic Themes. We note that only one Analytic Theme (Reflection on How Identities Affect Experiences) did not have subthemes so that the subtheme and analytic theme are the same in this case.
As an example, Table LABEL:tab:examples shows that for the first Analytic Theme “Pathways to Majoring in Physics”, one subtheme was “Family’s role in providing support”. The general codes for this were: Providing academic support, Providing encouragement to pursue science, and Providing exposure to science. Depending upon the narratives of the women, there were slight variations in their fine grain codes under this general code. We note that in the excerpts, we removed most of the “ums” and “likes” to make it easier to follow their point of view. We also note that occasionally some women elaborated more than others regarding certain instances, so some voices may come across as more emphatic pertaining to a particular issue than others.
III.4 Analytic Themes
@@
IV Findings and Discussion
Below, we present the findings from undergraduate women in physics at this institutions. The analytic themes are presented following their subthemes and general codes  [^2] General codes are not always bolded in the following sections because they were intertwined in the data. (bolded).
IV.1 Pathway to Majoring in Physics